OpenStudy (anonymous):

When 3 coins are tossed, the probability of getting 3 tails is 1/8 and the probability of 3 heads is also 1/8. The probability of 2 tails and 1 head is 3/8 and the probability of 2 heads and 1 tail is also 3/8. What is the probability of repeatedly tossing a group of three coins and getting 2 tails and 1 head 5 times in succession?

OpenStudy (anonymous):

(3/8)^5 = 243/32768 = 0.00741577148